OpenStudy (anonymous):

What is the end behavior of the function f(x) = −x^3 + 2x^2 + 4x + 5?

OpenStudy (anonymous):

Up on the left, up on the right Up on the left, down on the right Down on the left, up on the right Down on the left, down on the right

OpenStudy (mathmale):

Seriously, this question boils down to "What is the end behavior of f(x) = −x^3?" If you're not sure of what to do, graph g(x)=x^3. Then graph h(x)=-x^3.

OpenStudy (anonymous):

Thanks!